Circular No. 7198

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 H1 (LEE) H. Weaver, Johns Hopkins University; W. Vacca and N. Biver, Univer-

sity of Hawaii; S. Kim, Kyunghee University; D. Bockelée-Morvan and J. Crovisier, Observatoire de Paris-Meudon; T. Brooke, Jet Propulsion Laboratory; G. Chin, Goddard Space Flight Center, NASA; and J. Davies, Joint Astronomy Centre, report: “A high spectral resolution (R ∼ 10 000) observation made on May 20.3 UT with CSHELL at the Infrared Telescope Facility provided a clear detection of the ^{r}Q_{2} (F = 5.9 × 10^{−18} W m^{−2} in a 2′′ × 4′′ aperture) and ^{r}Q_{3} (F = 8.7 × 10^{−18} W m^{−2}) subbands of the ethane ν_{7} band near 3.35 μm, from which we estimate an ethane production rate of Q(C_{2}H_{6}) ∼ 0.7–1.4 × 10^{27} s^{−1}, assuming T (rot) = 50 K. The methane P2 lines were possibly detected at the 3σ level (F = 1.9 × 10^{−18} W m^{−2}) in the same spectral region, yielding a methane production rate of Q(CH_{4}) ∼ 4.5 × 10^{26} s^{−1}, assuming T (rot) = 50 K.”

SUPERNOVA 1999cg IN ANONYMOUS GALAXY A. Gal-Yam and D. Maoz, Tel Aviv University, and P. Guhathakurta,

University of California at Santa Cruz, write: “A spectrum of SN 1999cg (cf. IAUC 7181), obtained with the Keck-II telescope (+ LRIS spectrograph, range 380–880 nm) on May 30 by C. Sorenson and B. Schaefer, reveals that the supernova is of type Ia, ∼ 6 weeks past maximum brightness. The derived redshift of 0.14 is consistent with that of Abell 1607.”

SUPERNOVA 1999cl IN NGC 4501 P. Garnavich, S. Jha, R. Kirshner, and P. Challis, Harvard-Smithson-

ian Center for Astrophysics, report: “CCD imaging of SN 1999cl was obtained by A. Szentgyorgyi with the Fred L. Whipple Observatory 1.2-m telescope. Preliminary photometric reduction yields the following magnitudes: June 6.2 UT, V = 14.35, B − V = +1.01; 9.2, 14.03, +1.00 (±0.03 mag). The supernova brightness increased by 0.3 mag in three days, indicating that it is still well before maximum light. The color is consistent with 3 mag of visual extinction, as suggested by its spectrum (cf. IAUC 7190). Assuming A_{V} = 3.0 mag and that SN 1999cl is of normal luminosity, the distance to NGC 4501 (= M88) is < 12 Mpc, which is significantly smaller than the Tully-Fisher distance of 17 Mpc (Schoniger and Sofue 1997, A.Ap. 323, 14). Thus, the supernova may be overluminous, the extinction law may be abnormal, or the distance may have been overestimated. Observations of SN 1999cl, particularly in the infrared, are encouraged.”

1999 June 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green